# A Spectral Sequence for Equidimensional Actions of Compact Lie Groups

### v1

@@ 3.3. Construction and the First Page of the Spectral Sequence
But since $N_{0}$ is connected the isometries of the orbits have to be represented by elements from $N_{0}$ for non-generic orbits as well (by approaching the transition function on a non-generic orbit with ones on nearby generic orbits). Since such maps act trivially on the cohomology of generic orbits (by a homotopy argument), we can easilly glue such subsets via a version of the Mayer-Vietoris Theorem for $E^{p,q}_{0}=\{\Omega_{G}^{p,q}(V),d^{0,1}\}$:
@@
Let $(M,g)$ be a compact connected Riemannian manifold with an equidimensional action of $G$. Moreover, let $U$ and $V$ be open saturated subsets. Then the following short sequence of chain complexes is exact:
@@ 3.4. The second page of the spectral sequence
To compute the second page $E^{p,q}_{2}(M)$ it suffices to show that locally $d^{1,0}$ vanishes on elements from $\Omega^{\bullet}(\mathfrak{g},\mathfrak{h})^{(H\cap N_{0})/\penalty 50H_{0}}$, treated locally as a subspace of invariant forms via the splitting of $TM_{2}$. More precisely, let $V$ be a neighbourhood of any point $x\in M_{2}$ admitting a slice $D$ through $x$ perpendicular to $Gx$ and consider the quotient map $\pi:G\times D\to G\times D/\penalty 50G_{x}\cong V$. We consider an invariant Riemannian metric $g$ such as in the previous subsections. We can naturally identify $\Omega^{\bullet}(\mathfrak{g},\mathfrak{h})^{(H\cap N_{0})/\penalty 50H_{0}}$ with a subspace of $\Omega_{G}^{\bullet}(G\times D)$ using the orthogonal complement of the vertical bundle (which we denote by $T\tilde{\mathcal{F}}_{2}$ as in the previous section) for $G\times D$. We see that with the above identification we have $\Omega^{\bullet}(\mathfrak{g},\mathfrak{h})^{(H\cap N_{0})/\penalty 50H_{0}}\cap\pi^{*}(\Omega^{\bullet}_{G}(V))=\Omega^{\bullet}(\mathfrak{g},\mathfrak{h})^{G_{x}/\penalty 50G_{x0}}$, which due to $\pi^{*}$ being fiberwise monomorphic can be identified with a subspace of $\Omega^{\bullet}_{G}(V)$. Hence, our goal is to show that $d^{1,0}$ vanishes on this subspace.
@@
Let us use this frame to compute $d^{1,0}\tilde{\alpha}$ for $\tilde{\alpha}\in\Omega^{\bullet}(\mathfrak{g},\mathfrak{h})^{(H\cap N_{0})/\penalty 50H_{0}}$ (the lift of $\alpha$ to $G\times D$). For this we write for $Y\in\{Y_{1},...,Y_{q}\}$:
@@
and consider the terms on the right at a lift $\tilde{x}$ of $x$. The first term vanishes since by our choice of $X_{i}$ the function arising from evaluating $\tilde{\alpha}$ on these vector fields is constant. The second and fourth term vanish since they include the evaluation of $\tilde{\alpha}$ on a perpendicular vector field $Y$. Similarly, the third term has to be zero as well since each $Y,X_{i_{l}}]$ is perpendicular (since by Remark 3.2 the lifted metric is bi-invariant on $T(G\times D)|_{G\tilde{x}}$ with respect to the left and right multiplications in $G$). Consequently, $d^{1,0}\tilde{\alpha}=0$ at $\tilde{x}$ which implies that $d^{1,0}\alpha=0$ at $x$ (since the map induced on forms by the quotient $G\times D\to G/\penalty 50(H\cap N_{0})\times D\cong V$ is monomorphic). Hence, we have proved that $d^{1,0}$ vanishes locally on elements from $\Omega^{\bullet}(\mathfrak{g},\mathfrak{h})^{(H\cap N_{0})/\penalty 50H_{0}}$ at any point $x\in U$. To pass to general orbits it suffices to consider the Slice Theorem for a non-generic orbit and approach that orbit by points in $U$.
This finishes the proof of Theorem 1.1 since it implies that $d^{1,0}$ on the first page doesn’t affect $H^{\bullet}(\mathfrak{g},\mathfrak{h})$. To see this note that for $M_{2}$ classes in $H^{\bullet}(\mathfrak{g},\mathfrak{h})$ can be locally (in the forms on a neighbourhood admitting a slice) represented by elements in $\Omega^{\bullet}(\mathfrak{g},\mathfrak{h})^{N_{0}}\subset\Omega^{\bullet}(\mathfrak{g},\mathfrak{h})^{(H\cap N_{0})/\penalty 50H_{0}}$. Since the transition functions between neighbourhoods admitting slices preserve orthogonal complements and restricted to each $G$ orbit act by an element of $N_{0}$ these forms can be glued to global representatives of clasess in $\Omega^{0}(M_{2}/\penalty 50\mathcal{F}_{2},H^{\bullet}(\mathfrak{g},\mathfrak{h})^{N_{0}})\cong\Omega^{0}(M_{2}/\penalty 50\mathcal{F}_{2},H^{\bullet}(\mathfrak{g},\mathfrak{h}))$ on which $d^{1,0}$ vanishes. Consequently, passing to cohomology (with respect to $d^{1,0}$) will leave the coefficients (identified with the classes represented by the above constructed basic 0-forms) intact and hence the second page is just:

### v2

@@ 3.3. Construction and the First Page of the Spectral Sequence
But since $N_{0}$ is connected the isometries of the orbits have to be represented by elements from $N_{0}$ for non-generic orbits as well (by approaching the transition function on a non-generic orbit with ones on nearby generic orbits). Since such maps act trivially on the cohomology of orbits (by a homotopy argument for $H^{\bullet}_{dR}(G/\penalty 50H_{0})\cong H^{\bullet}(\mathfrak{g},\mathfrak{h})$), we can easilly glue such subsets via a version of the Mayer-Vietoris Theorem for $E^{p,q}_{0}=\{\Omega_{G}^{p,q}(V),d^{0,1}\}$:
@@
Let $(M,g)$ be a compact connected Riemannian manifold with an equidimensional action of $G$. Moreover, let $U$ and $V$ be open saturated subsets. Then the following short sequence of cochain complexes is exact:
@@ 3.4. The second page of the spectral sequence
To compute the second page $E^{p,q}_{2}(M)$ we will first show that locally $d^{1,0}$ vanishes on elements from $\Omega^{\bullet}(\mathfrak{g},\mathfrak{h})^{N_{0}}$, treated locally as a subspace of invariant forms via the splitting of $TM_{2}$. More precisely, let $V$ be a neighbourhood of any point $x\in M^{H_{0}}_{2}$ admitting a slice $D$ through $x$ perpendicular to $Gx$ and consider the quotient map $\pi:G\times D\to G\times D/\penalty 50G_{x}\cong V$. We consider an invariant Riemannian metric $g$ such as in the previous subsections. We can naturally identify $\Omega^{\bullet}(\mathfrak{g},\mathfrak{h})^{N_{0}}$ with a subspace of $\Omega_{G}^{\bullet}(G\times D/\penalty 50\mathcal{H})\subset\Omega_{G}^{\bullet}(G\times D)$ using the orthogonal complement of $T\tilde{\mathcal{F}_{2}}/\penalty 50T\mathcal{H}$ in $N\mathcal{H}$ (where $\tilde{\mathcal{F}_{2}}$, $\mathcal{H}$ and the metric on $N\mathcal{H}$ are as in the previous section). Note that all such forms descend to $\Omega^{\bullet}_{G}(V)$ since $G_{x}\subset N_{0}$ and hence $\Omega^{\bullet}(\mathfrak{g},\mathfrak{h})^{N_{0}}$ can be identified with a subspace of $\Omega^{\bullet}_{G}(V)$. Moreover, note that this space does not depend on the choice of slice through $x$ since if we repeat the construction for a different choice it would simply change each orbit by a right multiplication by some element of $N_{0}$ (possibly different for each orbit). Since, regardless of the choice of slice the descended forms vanish on the orthogonal complement of $T\mathcal{F}_{2}|_{V}$ and restricted to the orbit are $N_{0}$ invariant this shows that indeed this space does not depend on the chosen slice through $x$. Our goal is to show that $d^{1,0}$ vanishes on these descended forms.
@@
Let us use this frame to compute $d^{1,0}\tilde{\alpha}$ for $\tilde{\alpha}\in\Omega^{\bullet}(\mathfrak{g},\mathfrak{h})^{N_{0}}$ (the lift of $\alpha$ to $G\times D$). For this we write for $Y\in\{Y_{1},...,Y_{q}\}$:
@@
and consider the terms on the right at a lift $\tilde{x}$ of $x$. The first term vanishes since by our choice of $X_{i}$ the function arising from evaluating $\tilde{\alpha}$ on these vector fields is constant. The second and fourth term vanish since they include the evaluation of $\tilde{\alpha}$ on a perpendicular vector field $Y$. Similarly, the third term has to be zero as well since each $Y,X_{i_{l}}]$ is perpendicular (since by Remark 3.2 the lifted metric is bi-invariant on $T(G\times D)|_{G\tilde{x}}$ with respect to the left and right multiplications in $G$). Consequently, $d^{1,0}\tilde{\alpha}=0$ at $\tilde{x}$ which implies that $d^{1,0}\alpha=0$ at $x$ (since the map induced on forms by the quotient $G\times D\to G/\penalty 50(H\cap N_{0})\times D\cong V$ is pointwise monomorphic) and consequently $d^{1,0}\alpha=0$ at any point in $Gx$. To show that this is the case for other orbits through $y\in V$ we repeat this computation for a slice which is perpendicular to the chosen orbit (we use the fact that the space $\Omega^{\bullet}(\mathfrak{g},\mathfrak{h})^{N_{0}}$ does not depend on the choice of slice) which can be done since a connected component of the restriction of this slice to an appropriate saturated tubular neighbourhood of the chosen orbit is a slice in its own right. Hence, we have proved that $d^{1,0}$ vanishes locally on elements from $\Omega^{\bullet}(\mathfrak{g},\mathfrak{h})^{N_{0}}$ at any point $x\in U$. To pass to general orbits it suffices to consider the Slice Theorem for a non-generic orbit and approach that orbit by points in $U$.
This finishes the proof of Theorem 1.1 since it implies that $d^{1,0}$ on the first page doesn’t affect $H^{\bullet}(\mathfrak{g},\mathfrak{h})$. To see this note that for $M_{2}$ classes in $H^{\bullet}(\mathfrak{g},\mathfrak{h})$ can be locally (in the forms on a neighbourhood admitting a slice) represented by elements in $\Omega^{\bullet}(\mathfrak{g},\mathfrak{h})^{N_{0}}$. Since the transition functions between neighbourhoods admitting slices preserve orthogonal complements and restricted to each $G$ orbit act by an element of $N_{0}$ these forms can be glued to global representatives of clasess in $\Omega^{0}(M_{2}/\penalty 50\mathcal{F}_{2},H^{\bullet}(\mathfrak{g},\mathfrak{h})^{N_{0}})\cong\Omega^{0}(M_{2}/\penalty 50\mathcal{F}_{2},H^{\bullet}(\mathfrak{g},\mathfrak{h}))$ on which $d^{1,0}$ vanishes. Consequently, passing to cohomology (with respect to $d^{1,0}$) will leave the coefficients (identified with the classes represented by the above constructed basic 0-forms) intact and hence the second page is just: